OpenStudy (deaththekid975):

Maddie's bowling score was 230 last week, but only 190 this week. What was the percent decrease in Maddie's score? A. 1.74% B. 2.11% C. 17.4% D. 21.1%

OpenStudy (welshfella):

decrease = 230 - 190 divide this by last weeks score then multiply by 100.

OpenStudy (smartnerd1111):

so you do 230-190=40 and then do 40/190 to get your answer

OpenStudy (deaththekid975):

ok

OpenStudy (smartnerd1111):

then you multiply that answer by 100

OpenStudy (welshfella):

No - you divide by LAST weeks score The decrease is based on this.

OpenStudy (deaththekid975):

0.210

OpenStudy (smartnerd1111):

so you do 230-190=40 and then do 40/190 then multiply that answer by 100

OpenStudy (smartnerd1111):

@deaththekid975 multiply 0.210 by 100

OpenStudy (deaththekid975):

21.05

OpenStudy (smartnerd1111):

or 21.1

OpenStudy (welshfella):

NO that is incorrect last wekks score was 230 correct is (40/230) * 100

OpenStudy (smartnerd1111):

ohhhhhhhhhhhhhh

OpenStudy (aveline):

I second welshfella. @smartnerd1111 probably misread the question (like me at first xD). It's the percent decrease not increase

OpenStudy (smartnerd1111):

i misread it is 40/230 * 100

OpenStudy (deaththekid975):

17.39

OpenStudy (smartnerd1111):

or 17.4%

OpenStudy (deaththekid975):

ok! thanks